Q.For the quadratic equation 3xsquare + px + 3 + 0 , p>0.If one of the roots is square of the other,then p=
a) 1/3        b) 1        c) 3        d) none of these

Grade:12

4 Answers

Chetan Mandayam Nayakar
11 years ago

let roots be q,q2, product of roots= 3/3 =1, q3=1, q=1, sum of roots= -p/3

Ram Tiwari
11 years ago

sir i also tried the same thing
by that p is coming out to be -6
but the answer given is option (c)

shubham bhatia
11 years ago

since eqn says one root is square of other so if one root is q then other root is q square therfore qcube from product of roots is1 therefore q=1 or q=w(omega)complex root q=1 is rejected as p =-6 which is not possible therfore q=w and 1+w+wsquare=0 which is property of complex no since w is root there 3wsquare +pw +3=0 there p=3

Ram Tiwari
11 years ago

thank you sir